OpenStudy (czarluc):

25x^2-36y^2=1 How to turn this equation to its standard form. it is a hyperbola

OpenStudy (anonymous):

What if you divided by 900 throughout?

OpenStudy (michele_laino):

please note that I can re-write your equation as follows: \[\frac{{{x^2}}}{{{{\left( {1/5} \right)}^2}}} - \frac{{{y^2}}}{{{{\left( {1/6} \right)}^2}}} = 1\]

OpenStudy (michele_laino):

hint: \[\frac{{{x^2}}}{{\frac{1}{{25}}}} - \frac{{{y^2}}}{{\frac{1}{{36}}}} = 1\]

OpenStudy (michele_laino):

please keep in mind that: \[\frac{1}{{25}} = {\left( {\frac{1}{5}} \right)^2}\quad \frac{1}{{36}} = {\left( {\frac{1}{6}} \right)^2}\]

OpenStudy (anonymous):

Use the formula that you find in the book then plug in the numbers to get the right answers. always relate to the book